OpenStudy (anonymous):

What is the simplified form of 5/y^2 - 4/3y

OpenStudy (anonymous):

that is the simplified form only 4y/3. You had the first part correct though.

OpenStudy (anonymous):

@KE8717504 No, that's what they put.

OpenStudy (anonymous):

The options are 5y - 12/3y^2 4y - 15/ 3y^2 12 - 5y / 3y^2 15 - 4y / 3y^2

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

\[\Large \frac{5}{y^2} - \frac{4}{3y}\] \[\Large \frac{3*5}{3*y^2} - \frac{4}{3y}\] \[\Large \frac{15}{3y^2} - \frac{4}{3y}\] \[\Large \frac{15}{3y^2} - \frac{4y}{3y*y}\] \[\Large \frac{15}{3y^2} - \frac{4y}{3y^2}\] \[\Large \frac{15-4y}{3y^2}\] ------------------------------------------------------- So \[\Large \frac{5}{y^2} - \frac{4}{3y}\] simplifies to \[\Large \frac{15-4y}{3y^2}\]

OpenStudy (anonymous):

Ahhhh you're the best!!!!!

jimthompson5910 (jim_thompson5910):

Thanks for the compliment. I'm glad to be of help to you.